# Linear cost and exponentially convergent approximation of Gaussian Matérn processes

### v1

@@ Abstract
The computational cost for inference and prediction of statistical models based on Gaussian processes with Matérn covariance functions scales cubicly with the number of observations, limiting their applicability to large data sets. The cost can be reduced in certain special cases, but there are currently no generally applicable exact methods with linear cost. Several approximate methods have been introduced to reduce the cost, but most of these lack theoretical guarantees for the accuracy. We consider Gaussian processes on bounded intervals with Matérn covariance functions and for the first time develop a generally applicable method with linear cost and with a covariance error that decreases exponentially fast in the order $m$ of the proposed approximation. The method is based on an optimal rational approximation of the spectral density and results in an approximation that can be represented as a sum of $m$ independent Gaussian Markov processes, which facilitates easy usage in general software for statistical inference, enabling its efficient implementation in general statistical inference software packages. Besides the theoretical justifications, we demonstrate the accuracy empirically through carefully designed simulation studies which show that the method outperforms all state-of-the-art alternatives in terms of accuracy for a fixed computational cost in statistical tasks such as Gaussian process regression.
keywords:
Gaussian process, Gaussian Markov random field, inference, prediction
@@ 1 Introduction
Gaussian stochastic processes with Matérn covariance functions (Matérn, 1960) are important models in statistics and machine learning (Porcu et al., 2023), and in particular in areas such as spatial statistics (Stein, 1999; Lindgren et al., 2022), computer experiments (Santner et al., 2003; Gramacy, 2020) and Bayesian optimization (Srinivas et al., 2009). A Gaussian process $u$ on $\mathbb{R}$ has a Matérn covariance function if Cov$(u(s),u(t))=r(s,t;\nu,\kappa,\sigma^{2})$ with
$r(s,t;\nu,\kappa,\sigma^{2})=\varrho(|s-t|;\nu,\kappa,\sigma^{2}),\quad\varrho(h;\nu,\kappa,\sigma^{2})=\frac{\sigma^{2}}{2^{\nu-1}\Gamma{(\nu)}}(\kappa h)^{\nu}K_{\nu}(\kappa h),$ (1)
@@
where $K_{\nu}(\cdot)$ is a modified Bessel function of the second kind and $\Gamma(\cdot)$ denotes the gamma function. The three parameters $\kappa,\sigma^{2},\nu>0$, determine the practical correlation range, variance and smoothness of the process, respectively.
@@
However, the approach of defining Gaussian processes via the covariance function incurs a significant computational cost for inference and prediction, mainly due to the requirement of factorising dense covariance matrices. To address this so called “Big n” problem, referring to the $\mathcal{O}(n^{3})$ computational cost required for problems with $n$ observations, several methods have been developed in recent years to reduce the cost. In this work, we focus on the case where the data is observed on an interval $I\subset R$. In one spatial dimension, the processes have Markov properties if $\alpha=\nu+\nicefrac{{1}}{{2}}\in\mathbb{N}$, and this has been used to derive exact methods with a $\mathcal{O}(n)$ cost for these particular cases, see for example the Kernel packet method Chen et al. (2022) or the state-space methods of Hartikainen & Särkkä (2010); Särkkä & Hartikainen (2012); Särkkä et al. (2013). If the data is evenly spaced in $\mathbb{R}$, Toeplitz methods Wood & Chan (1994) can also be used to reduce the computational cost to $\mathcal{O}(n\log^{2}n)$ Ling (2019); Ling & Lysy (2022). Outside these particular cases, there are currently no exact methods that can reduce the computational cost, and one instead needs to rely on approximations. One popular method is the random Fourier features approach of Rahimi & Recht (2007) which gives a cost $\mathcal{O}(m^{2}+mn)$ when using $m$ features, and an accuracy of $\mathcal{O}(m^{-\nicefrac{{1}}{{2}}})$. This is one of the few methods that have theoretical guarantees for how quickly the approximation error decreases as the order $m$ increases. Unfortunately, the theoretical rate is low and as we will later see, the method provides poor approximations for Matérn processes on $\mathbb{R}$.
@@
Another widely used method is the SPDE approach of (Lindgren et al., 2011), where the Matérn process is represented as a solution to a stochastic differential equation (SDE) which is approximated via a finite element method (FEM) approximation. This was originally proposed for the case $\nu-\nicefrac{{1}}{{2}}\in\mathbb{N}_{0}$ and was later extended by Bolin et al. (2020, 2024b) to general $\nu>0$, where the authors also derived explicit rates of convergence of the approximation in terms of the finite element mesh width. The approach is computationally efficient, but has the disadvantage that the SDE has certain boundary conditions which makes the approximation converge to a non-stationary covariance which is only similar to the Matérn covariance away from the boundary of the computational domain. The rate of convergence is better than that of the random Fourier features method, but does not decrease exponentially fast. Other approaches, which are generally applicable, but without theoretical rates of convergence of the covariance function approximation are covariance tapering Furrer et al. (2006), Vecchia approximation Vecchia (1988); Gramacy & Apley (2015); Datta et al. (2016), low rank methods (Higdon, 2002; Cressie & Johannesson, 2008), and multiresolution approximations (Nychka et al., 2015). The state space methods have also been extended to general $\nu>0$ in Karvonen & Särkkä (2016) and Tronarp et al. (2018) through spectral transformation methods and the approximation of Matérn kernel by a finite scale mixture of squared exponential kernels, respectively. However, the accuracy the resulting approximated covariance function has been demonstrated only through numerical experiments and no theoretical analysis of the rate has been provided.
In this work, we develop a new method for Gaussian Matérn processes on intervals, which has at most $\mathcal{O}(nm^{3}\lceil{\alpha\rceil^{3}})$ computational cost when used for statistical inference, prediction and sampling, where $m$ is the order of the approximation. We prove in Section 2 that the error of the covariance function converges exponentially fast in $m$, which in practice means that $m$ can be chosen very low. Specifically, the error is $O\bigl{(}\exp(-2\pi\sqrt{\{\alpha\}m})\bigr{)}$ where $\{\alpha\}$ is the decimal part of $\alpha\notin\mathbb{N}$. If $\alpha\in\mathbb{N}$, the method is exact and has a cost of $\mathcal{O}(n\lceil{\alpha\rceil^{3}})$.
@@
The approach is based on a rational approximation of the spectral density, which thus is similar in spirit to (Karvonen & Särkkä, 2016; Roininen et al., 2018). The difference is, however, that we can derive theoretical guarantees for the error. Because the method is based on a rational approximation, it can be used in combination with state-space methods for efficient inference. We, however, derive in Section 3 a direct representation of the approximated process as a sum of Gaussian Markov random fields with sparse precision (inverse covariance) matrices, which in fact are band matrices. This representation has several advantages, and perhaps the most important is that it can directly be incorporated in general software for Bayesian inference, such as R-INLA (Lindgren & Rue, 2015). Another important feature of the proposed method is that the linear cost is applicable to working with the process and its derivatives jointly (and not only the process). This is useful in several applications that arise in the natural sciences where the observation of the derivatives is available, see for instance Solak et al. (2002); Padidar et al. (2021); Roos et al. (2021); Yang et al. (2018).
@@
To validate the effectiveness and accuracy of the method, we compare it in Section 4 to the covariance-based rational approximation method of Bolin et al. (2024b) as well as the Vecchia approximations (Datta et al., 2016), which recently were shown to be the best two methods for approximating Gaussian Matérn fields for geostatistical applications (Hong et al., 2023). We also compare with the state-space approach of Karvonen & Särkkä (2016), the random Fourier features method of Rahimi & Recht (2007) and the covariance tapering approach of Nychka et al. (2015). We show that the method outperforms the alternatives in terms of accuracy for a fixed cost when used for Gaussian process regression. The comparison also includes a principal component analysis (PCA) (Wang, 2008) approach, which serves as an “optimal” low-rank method. As the proposed method outperforms this, it means that it also would outperform any other low-rank method, such as fixed rank kriging (Cressie & Johannesson, 2008) or process convolutions (Higdon, 2002). Extensions of the method beyond stationary Matérn processes on intervals, as well as concluding remarks, are given in Section 5. The proposed method is implemented in the R package rSPDE (Bolin & Simas, 2023) available on CRAN, and all code for the comparisons, as well as a Shiny application with further results can be found in https://github.com/vpnsctl/MarkovApproxMatern/. Proofs and technical details are provided in two technical appendices.
@@ 2 Exponentially convergent rational approximation
Let $u$ be a centered Gaussian Process on an interval $I\subset\mathbb{R}$ with covariance function (1) and $\alpha=\nu+\nicefrac{{1}}{{2}}\notin\mathbb{N}$ (which is the case for which no exact and efficient methods exist). This process has spectral density $f^{\alpha}(w)=A\sigma^{2}(\kappa^{2}+w^{2})^{-\alpha},$ where $A=\frac{1}{2\pi}\Gamma(\alpha)\sqrt{4\pi}\kappa^{2\nu}\Gamma(\nu)^{-1}$ Lindgren (2012). We define a rational approximation of the process $u$ as a Gaussian process with spectral density
$f_{m}^{\alpha}(w)=A\kappa^{-2\alpha}\frac{\sigma^{2}}{(1+\kappa^{-2}w^{2})^{\lfloor\alpha\rfloor}}\frac{P_{m}(1+\kappa^{-2}w^{2})}{Q_{m}(1+\kappa^{-2}w^{2})},$ (2)
@@
where $P_{m}$ and $Q_{m}$ are polynomials derived from the optimal rational approximation of order $m$ for the real-valued function $f(x)=x^{\{\alpha\}}$ on the interval $[0,1]$, with respect to the supremum norm. That is,
$x^{\{\alpha\}}\approx\frac{P_{m}(x)}{Q_{m}(x)}=\frac{\sum_{i=0}^{m}a_{i}x^{i}}{\sum_{i=0}^{m}b_{i}x^{i}},$
@@
where this approximation is the best with respect to the supremum norm on $[0,1]$. The coefficients $\{a_{i}\}_{i=0}^{m}$ and $\{b_{i}\}_{i=0}^{m}$ in the rational approximation can be obtained via the second Remez algorithm (Remez, 1934) or by using the recent, and more stable, BRASIL algorithm (Hofreither, 2021). An additional option to compute the coefficients is the Clenshaw-Lord Chebyshev-Padé algorithm (Baker & Morris, 1996) which provides a robust “near best” rational approximation.
By performing a partial fraction decomposition of the rational function $P_{m}(x)/Q_{m}(x)$ in (2), we can write the spectral density of the approximation as
@@
$f_{m}^{\alpha}(w)$ $=A\sigma^{2}\kappa^{-2\alpha}\left[\frac{k}{(1+\kappa^{-2}w^{2})^{\lfloor\alpha\rfloor}}+\sum_{i=1}^{m}r_{i}\frac{1}{(1+\kappa^{-2}w^{2})^{\lfloor\alpha\rfloor}(1+\kappa^{-2}w^{2}-p_{i})}\right]$ (3)
$=:\left[f_{m,0}^{\alpha}(w)+\sum_{i=1}^{m}f_{m,i}^{\alpha}(w)\right],$

### v2

@@ Abstract
The computational cost for inference and prediction of statistical models based on Gaussian processes with Matérn covariance functions scales cubically with the number of observations, limiting their applicability to large data sets. The cost can be reduced in certain special cases, but there are no generally applicable exact methods with linear cost. Several approximate methods have been introduced to reduce the cost, but most lack theoretical guarantees for accuracy. We consider Gaussian processes on bounded intervals with Matérn covariance functions and, for the first time, develop a generally applicable method with linear cost and a covariance error that decreases exponentially fast in the order $m$ of the proposed approximation. The method is based on an optimal rational approximation of the spectral density and results in an approximation that can be represented as a sum of $m$ independent Gaussian Markov processes, facilitating usage in general software for statistical inference. Besides theoretical justifications, we demonstrate accuracy empirically through carefully designed simulation studies, which show that the method outperforms state-of-the-art alternatives in accuracy for fixed computational cost in tasks like Gaussian process regression.
Keywords: Gaussian process, Gaussian Markov random field, inference, prediction
@@ 1 Introduction
Gaussian stochastic processes with Matérn covariance functions (Matérn, 1960) are important models in statistics and machine learning (Porcu et al., 2023), and in particular in areas such as spatial statistics (Stein, 1999; Lindgren et al., 2022), computer experiments (Santner et al., 2003; Gramacy, 2020) and Bayesian optimization (Srinivas et al., 2009). Although the Matérn covariance often is used for spatial data, it is also commonly used for temporal data, in particular in areas such as functional data analysis (Yang et al., 2016), longitudinal data analysis (Asar et al., 2020), and growth rate modeling (Swain et al., 2016). A Gaussian process $u$ on $\mathbb{R}$ has a Matérn covariance function if
$\operatorname{Cov}(u(s),u(t))=\varrho(|s-t|;\nu,\kappa,\sigma^{2}),\quad\varrho(h;\nu,\kappa,\sigma^{2})=\frac{\sigma^{2}}{2^{\nu-1}\Gamma{(\nu)}}(\kappa h)^{\nu}K_{\nu}(\kappa h),$ (1)
@@
where $K_{\nu}(\cdot)$ is a modified Bessel function of the second kind of order $\nu$, and $\Gamma(\cdot)$ denotes the gamma function. The three parameters $\kappa,\sigma^{2},\nu>0$, determine the practical correlation range, variance and smoothness of the process, respectively. We introduce the practical correlation range $\rho=\sqrt{8\nu}/\kappa$ as the distance at which the correlation is approximately 0.1, and a new smoothness parameter $\alpha=\nu+\nicefrac{{1}}{{2}}$, which will be used throughout this text.
@@
However, the approach of defining Gaussian processes via the covariance function incurs a significant computational cost for inference and prediction, mainly due to the requirement of factorising dense covariance matrices. To address this so called “Big n” problem, referring to the $\mathcal{O}(n^{3})$ computational cost required for problems with $n$ observations, several methods have been developed in recent years to reduce the cost. In this work, we consider the case where the data is observed on an interval $I\subset\mathbb{R}$. The processes have Markov properties when $\alpha\in\mathbb{N}$, and this has been used to derive exact methods with a $\mathcal{O}(n)$ cost for these particular cases, see for example the Kernel packet method (Chen et al., 2022) or the state-space methods of Hartikainen and Särkkä (2010); Särkkä and Hartikainen (2012); Särkkä et al. (2013). If the data is evenly spaced in $\mathbb{R}$, Toeplitz methods (Wood and Chan, 1994) can also be used to reduce the computational cost to $\mathcal{O}(n(\log n)^{2})$ (Ling, 2019; Ling and Lysy, 2022). Outside these particular cases, there are currently no exact methods that can reduce the computational cost, and one instead needs to rely on approximations. One popular method is the random Fourier features approach of Rahimi and Recht (2007) which gives a cost $\mathcal{O}(m^{2}+mn)$ when using $m$ features, and an accuracy of $\mathcal{O}(m^{-\nicefrac{{1}}{{2}}})$. This is one of the few methods that have theoretical guarantees for how quickly the approximation error decreases as the order $m$ increases. Unfortunately, the theoretical rate is low and as we will later see, the method provides poor approximations for Matérn processes on $\mathbb{R}$.
@@
Another widely used method is the SPDE approach of Lindgren et al. (2011), where the Matérn process is represented as a solution to a stochastic differential equation (SDE) which is approximated via a finite element method (FEM) approximation. This was originally proposed for the case $\nu-\nicefrac{{1}}{{2}}\in\mathbb{N}_{0}$ (i.e., $\alpha\in\mathbb{N}$) and was later extended by Bolin et al. (2020, 2024b) to general $\nu>0$, where the authors also derived explicit rates of convergence of the approximation in terms of the finite element mesh width. The approach is computationally efficient, but has the disadvantage that the SDE has certain boundary conditions which makes the approximation converge to a non-stationary covariance which is only similar to the Matérn covariance away from the boundary of the computational domain. The rate of convergence is better than that of the random Fourier features method, but does not decrease exponentially fast. Other approaches, which are generally applicable, but without theoretical rates of convergence of the covariance function approximation are covariance tapering (Furrer et al., 2006), Vecchia approximation (Vecchia, 1988; Gramacy and Apley, 2015; Datta et al., 2016), low rank methods (Higdon, 2002; Cressie and Johannesson, 2008), and multiresolution approximations (Nychka et al., 2015). The state space methods have also been extended to general $\nu>0$ in Karvonen and Särkkä (2016) and Tronarp et al. (2018) through spectral transformation methods and the approximation of the Matérn kernel by a finite scale mixture of squared exponential kernels, respectively. However, the accuracy the resulting approximated covariance function has been demonstrated only through numerical experiments and no theoretical analysis of the rate has been provided.
@@
In this work, we develop a new method for Gaussian Matérn processes on intervals, which has at most $\mathcal{O}(nm^{3}\lceil{\alpha\rceil^{3}})$ computational cost when used for statistical inference, prediction and sampling, where $m$ is the order of the approximation. We prove in Section 2 that the error of the covariance function converges exponentially fast in $m$, which in practice means that $m$ can be chosen very low. Specifically, the error is $O\bigl{(}\exp(-2\pi\sqrt{\{\alpha\}m})\bigr{)}$ where $\{\alpha\}$ is the fractional part of $\alpha\notin\mathbb{N}$. If $\alpha\in\mathbb{N}$, the method is exact and has a cost of $\mathcal{O}(n\lceil{\alpha\rceil^{3}})$. The approach is based on a rational approximation of the spectral density, which thus is similar in spirit to Karvonen and Särkkä (2016); Roininen et al. (2018). The difference is, however, that we have theoretical guarantees for the error. Because the method is based on a rational approximation, it can be used in combination with state-space methods for efficient inference. We, however, derive in Section 3 a direct representation of the approximated process as a sum of Gaussian Markov random fields with sparse precision (inverse covariance) matrices, which in fact are band matrices. This representation has several advantages, and perhaps the most important is that it can directly be incorporated in general software for Bayesian inference, such as R-INLA (Lindgren and Rue, 2015). Another important feature is that the linear cost is applicable to working with the process and its derivatives jointly. This is useful in several applications that arise in the natural sciences where observations of the derivatives are available (see, e.g. Solak et al., 2002; Padidar et al., 2021; Roos et al., 2021; Yang et al., 2018) or when derivatives are of direct interest (Swain et al., 2016).
@@
To validate the effectiveness and accuracy of the method, we compare it in Section 4 to the covariance-based rational approximation method of Bolin et al. (2024b) as well as the Vecchia approximations (Datta et al., 2016). These methods demonstrated superior performance among several alternatives for approximating Gaussian Matérn fields in specific geostatistical test problems, as shown in the comparative study by Hong et al. (2023). We also compare with the state-space approach of Karvonen and Särkkä (2016), the random Fourier features method of Rahimi and Recht (2007) and the covariance tapering approach of Nychka et al. (2015). We show that the method outperforms the alternatives in terms of accuracy for a fixed cost when used for Gaussian process regression. The comparison also includes a principal component analysis (PCA) (Wang, 2008) approach, which serves as an “optimal” low-rank method. As the proposed method outperforms this, it means that it also would outperform any other low-rank method, such as fixed rank kriging (Cressie and Johannesson, 2008) or process convolutions (Higdon, 2002). Extensions of the method beyond stationary Matérn processes on intervals, as well as concluding remarks, are given in Section 5. The proposed method is implemented in the R package rSPDE (Bolin and Simas, 2023) available on CRAN, and all code for the comparisons, as well as a Shiny application with further results can be found in https://github.com/vpnsctl/MarkovApproxMatern/. Proofs and technical details are provided in two technical appendices.
@@ 2 Exponentially convergent rational approximation
Let $u$ be a centered Gaussian Process on an interval $I\subset\mathbb{R}$ with covariance function (1) and $\alpha=\nu+\nicefrac{{1}}{{2}}\notin\mathbb{N}$ (which is the case for which no exact and efficient methods exist). This process has spectral density $f_{\alpha}(w)=A\sigma^{2}(\kappa^{2}+w^{2})^{-\alpha},$ where $A=\sqrt{2}\kappa^{2\nu}\Gamma(\nu+\nicefrac{{1}}{{2}})\Gamma(\nu)^{-1}$ (Lindgren, 2012). We define a rational approximation of the process $u$ as a Gaussian process with spectral density
$f_{m,\alpha}(w)=A\kappa^{-2\alpha}\frac{\sigma^{2}}{(1+\kappa^{-2}w^{2})^{\lfloor\alpha\rfloor}}\frac{P_{m}(1+\kappa^{-2}w^{2})}{Q_{m}(1+\kappa^{-2}w^{2})},$ (2)
@@
where $P_{m}(x)=\sum_{i=0}^{m}a_{i}x^{m-i}$ and $Q_{m}=\sum_{i=0}^{m}b_{i}x^{m-i}$ are polynomials derived from the optimal rational approximation of order $m$ for the real-valued function $f(x)=x^{\{\alpha\}}$ on the interval $[0,1]$, with respect to the supremum norm, where we recall that $\{\alpha\}$ is the fractional part of $\alpha$. More precisely, we use the approximation $x^{-\{\alpha\}}=\left(x^{-1}\right)^{\{\alpha\}}\approx P_{n}(x)/Q_{n}(x)$, where the coefficients $\{a_{i}\}_{i=0}^{m}$ and $\{b_{i}\}_{i=0}^{m}$ are such that the approximation
$x^{\{\alpha\}}\approx\frac{\sum_{i=0}^{m}a_{i}x^{i}}{\sum_{i=0}^{m}b_{i}x^{i}}$ (3)
@@
is the best with respect to the supremum norm on $[0,1]$. The coefficients $\{a_{i}\}_{i=0}^{m}$ and $\{b_{i}\}_{i=0}^{m}$ in the rational approximation are unique (Lorentz et al., 1996, Chapter 7.2) and can be obtained via the second Remez algorithm (Remez, 1934) or by using the recent, and more stable, BRASIL algorithm (Hofreither, 2021). Note that the polynomials $P_{m}(x)$ and $Q_{m}$ in (2) are written in terms of $x^{m-i}$ instead of $x^{i}$, using the coefficients from (3), because the rational approximation is applied to $(x^{-1})^{\{\alpha\}}$.
Figure 1: True and approximate spectral densities for $\nu=0.4$, $1.4$ and $2.4$, and for different rational approximation orders (top row), and corresponding absolute errors for the rational approximations (bottom row).
The true and approximate spectral densities for $\nu=0.4$, $1.4$ and $2.4$, along with their absolute errors, are shown in Figure 1 for different orders of rational approximation, where we obtained the coefficients using BRASIL algorithm. Here, we consider $\sigma=1$ and $\rho=2$, where $\rho=\nicefrac{{\sqrt{8\nu}}}{{\kappa}}$ represents the practical correlation range. We can observe that the approximation is very good even for low orders, with an almost perfect match already for order 2.
Let $r_{m,\alpha}$ be the corresponding covariance function obtained from the rational approximation of order $m$ of $f_{\alpha}$ given by (2), more precisely, let
[…]
An explicit expression for $r_{m,\alpha}(\cdot)$ will be obtained in the next section.
[…]
Theorem 1
Let $r_{\alpha}$ be the covariance function
$r_{\alpha}(s,t;\kappa,\sigma^{2})=\varrho(|s-t|;\alpha-\nicefrac{{1}}{{2}},\kappa,\sigma^{2}),\quad s,t,\in\mathbb{R},$ (5)
where $\varrho(\cdot)$ is the Matérn covariance function defined in (1) and $\alpha=\nu+\nicefrac{{1}}{{2}}$. Further, let $r_{m,\alpha}$ be the rational approximation given in (4). If $\alpha>1/2$, then
@@
$\|r_{m,\alpha}-r_{\alpha}\|_{L_{2}(I\times I)}\leq 2\sqrt{\pi}(b-a)A\sigma^{2}\kappa^{-2\alpha}\min\{1,M_{\lfloor\alpha\rfloor,\kappa}\}C_{\{\alpha\}}\mathbb{I}_{\alpha\notin\mathbb{N}}e^{-2\pi\sqrt{\{\alpha\}m}},$ (6)
where $\{\alpha\}$ denotes the fractional part of $\alpha$, $M_{n,\kappa}=\kappa\pi\Gamma(2n-1)/(4^{n-1}\Gamma(n)^{2}),$ $n\geq 1$ and $M_{0,\kappa}=\infty$. Additionally, $C_{\{\alpha\}}\in(0,\infty)$ is a constant that depends only on $\{\alpha\}$. Further, if $\alpha>1$, then a constant $K_{\alpha}\in(0,\infty)$ that only depends on $\alpha$ exists such that
$\|r_{m,\alpha}-r_{\alpha}\|_{C(I\times I)}\leq A\sigma^{2}\kappa^{-2\alpha}K_{\alpha}\mathbb{I}_{\alpha\notin\mathbb{N}}e^{-2\pi\sqrt{\{\alpha\}m}}.$ (7)
Remark 2
The constant $C_{\{\alpha\}}$ in (6) comes from the error of the rational approximation
$\sup_{x\in[0,1]}\left|x^{\{\alpha\}}-\frac{p_{m}(x)}{q_{m}(x)}\right|\leq C_{\{\alpha\}}e^{-2\pi\sqrt{\{\alpha\}m}},$
where $C_{\{\alpha\}}$ is independent of $m$. Further, for $\alpha\in(0,1)$, by Saff and Stahl (1995, Theorem 2), there is an approximate expression for this uniform error that is valid for large $m$:
$\sup_{x\in[0,1]}\left|x^{\{\alpha\}}-\frac{p_{m}(x)}{q_{m}(x)}\right|=4^{{\{\alpha\}}+1}|\sin(\pi{\{\alpha\}})|e^{-2\pi\sqrt{\{\alpha\}m}}(1+o(1)),\quad\text{as }m\to\infty.$
This expression provides an intuition on how the constant $C_{\{\alpha\}}$ behaves.
The theoretical error (6) using the approximate expression of $C_{\{\alpha\}}$ given in Remark 2 is shown in Figure 2 for $\rho=2$, $\sigma=1$, and different values of $\nu$ and $m$ on the interval $I=[0,50]$. We can observe that the errors do not decay monotonically as $\{\alpha\}$ increases, but instead, they have a more complex behavior with respect to $\alpha$. However, they decay monotonically as $m$ increases. One should note that the true error bounds may look a bit different compared to those in Figure 2 as they are based on using $4^{{\{\alpha\}}+1}|\sin(\pi{\{\alpha\}})|$ in place of $C_{\{\alpha\}}$. See Figure 3 for the actual covariance errors for this example.
Figure 2: Theoretical error bound from (6) in Theorem 1 using the approximation for $C_{\{\alpha\}}$ given in Remark 2.
3 Linear cost inference
Our goal now is to use the rational approximation in (2) to obtain a linear cost approximation of the covariance function of the Gaussian process $u$. This approximation will based on a partial fractions decomposition of the rational function in the approximate spectral density. Therefore, we first derive an important property of such partial fractions decompositions.
Proposition 3
Fix $\alpha\in(0,1)$ and $m\in\mathbb{N}$. Let the coefficients $\{a_{i}\}_{i=0}^{m}$ and $\{b_{i}\}_{i=0}^{m}$ be such that the best rational approximation of $x^{\alpha}$ on $[0,1]$ is given by (3). Further, let $P_{m}(x)=\sum_{i=0}^{m}a_{i}x^{m-i}$ and $Q_{m}=\sum_{i=0}^{m}b_{i}x^{m-i}$. Then, we have the following partial fractions decomposition of $P_{m}(x)/Q_{m}(x)$:
$\frac{P_{m}(x)}{Q_{m}(x)}=k+\sum_{i=1}^{m}\frac{c_{i}}{x-p_{i}},$
where $k,c_{i}>0$ and $p_{i}<0$ for $i=1,\ldots,m$.
Remark 4
Proposition 3 fills a theoretical gap left in Bolin et al. (2024b), where such a decomposition, along with the signs of $k$, $c_{i}$, and $p_{i}$ for $i=1,\ldots,m$, was verified numerically.
In view of Proposition 3, we can perform a partial fraction decomposition of the rational function $P_{m}(x)/Q_{m}(x)$ in (2) to obtain that the spectral density of the approximation is
$f_{m,\alpha}(w)$ $=A\sigma^{2}\kappa^{-2\alpha}\left[\frac{k}{(1+\kappa^{-2}w^{2})^{\lfloor\alpha\rfloor}}+\sum_{i=1}^{m}c_{i}\frac{1}{(1+\kappa^{-2}w^{2})^{\lfloor\alpha\rfloor}(1+\kappa^{-2}w^{2}-p_{i})}\right]$ (8)
$=:\left[f_{m,0,\alpha}(w)+\sum_{i=1}^{m}f_{m,i,\alpha}(w)\right],$
where $k,c_{i}>0$ and $p_{i}<0$ for $i=1,\ldots,m$, and $A=\sqrt{2}\kappa^{2\nu}\Gamma(\nu+\nicefrac{{1}}{{2}})\Gamma(\nu)^{-1}$. Thus, we obtain that $f_{m,\alpha}(\cdot)$ can be decomposed as a sum of valid spectral densities. Hence, we can write $r_{m,\alpha}(\cdot,\cdot)$, given in (4), as a sum of covariance functions.